OpenStudy (anonymous):

radical12 times radical 20?

OpenStudy (anonymous):

i assume square root hence answer is 4sqrt(15)

OpenStudy (anonymous):

\[\sqrt{a} * \sqrt{b} = \sqrt{ab}\]

OpenStudy (anonymous):

\[4\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{12}\times \sqrt{20}\]=\[4\sqrt{5}\]